OpenStudy (anonymous):

Nicolas has completed the experimental portion of the lab. Which of these should he do while he works on the analysis questions? work on a new experiment he thought of keep his safety goggles on talk to his friends who are still collecting data keep his gloves on

OpenStudy (anonymous):

@SuperflyTyGuy

OpenStudy (anonymous):

medal Plus fan

OpenStudy (superflytyguy):

I think it's talk to his friends who are still collecting data

OpenStudy (anonymous):

it was b

OpenStudy (superflytyguy):

Who wears goggles when you're done with the experiment?

OpenStudy (anonymous):

i think it is d

OpenStudy (anonymous):

Of the choices below, which unit of measurement would be most appropriate for measuring the length of a swimming pool? millimeters centimeters meters kilometers

OpenStudy (anonymous):

kilometers

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

meters

OpenStudy (anonymous):

Density is a measurable property of matter. How can the density of a substance be determined? Measure the mass and volume of the substance, then divide the volume by the mass. Measure the length, width, and height of the substance, then multiply these values together. Measure the mass and volume of the substance, then divide the mass by the volume. Measure the length and weight of the substance, then multiply the length by the weight.

OpenStudy (anonymous):

@hmann76

OpenStudy (anonymous):

@matlee

OpenStudy (anonymous):

Divide the mass by the volume: \(D=\frac {m}{V}\)

OpenStudy (anonymous):

who said u would have to me so mean about it

OpenStudy (anonymous):

Convert to millimeters: 28.8 meters

OpenStudy (superflytyguy):

0.0288

OpenStudy (anonymous):

\(28.8\times 10^3 = 28800mm\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

What is the equivalent of 13,245 seconds in hours, minutes, and seconds?